# Towards an Understanding of Context Utilization in Code Intelligence

### v1

@@ 7.1. Metric
Code Generation Python  (Tang et al., 2023; Agashe et al., 2019; Yu et al., 2024) (Liao et al., 2024; Xie et al., 2024; Li et al., 2024c; Hai et al., 2024) (Zhang et al., 2024b; Zheng et al., 2024c; Zhuo et al., 2024)  (Zheng et al., 2024c)  (Zan et al., 2023, 2022a) (Ma et al., 2024a) -  (Tang et al., 2023; Yu et al., 2024; Liao et al., 2024) (Li et al., 2024c; Hai et al., 2024; Zheng et al., 2024c; Zhuo et al., 2024) $\circ$ (Zheng et al., 2024c) $\circ$ (Tang et al., 2023; Agashe et al., 2019; Yu et al., 2024) (Liao et al., 2024; Xie et al., 2024; Li et al., 2024c) (Hai et al., 2024; Zheng et al., 2024c; Zhuo et al., 2024) $\circ$ (Tang et al., 2023; Agashe et al., 2019; Yu et al., 2024) (Liao et al., 2024; Xie et al., 2024; Li et al., 2024c) (Hai et al., 2024; Zheng et al., 2024c; Zhuo et al., 2024) $\circ$ (Tang et al., 2023; Agashe et al., 2019; Yu et al., 2024) (Liao et al., 2024; Xie et al., 2024; Li et al., 2024c) (Hai et al., 2024; Zheng et al., 2024c; Zhuo et al., 2024) $\circ$ (Tang et al., 2023; Yu et al., 2024; Liao et al., 2024) (Li et al., 2024c; Hai et al., 2024; Zheng et al., 2024c; Zhuo et al., 2024) $\circ$ (Tang et al., 2023; Yu et al., 2024; Liao et al., 2024) (Li et al., 2024c; Hai et al., 2024; Zheng et al., 2024c; Zhuo et al., 2024) - $\circ$ (Zan et al., 2023, 2022a; Tang et al., 2023) (Yu et al., 2024; Xie et al., 2024; Li et al., 2024c; Hai et al., 2024) (Zhang et al., 2024b; Zheng et al., 2024c; Zhuo et al., 2024) -
@@
Code Completion Python  (Pei et al., 2023; Ding et al., 2023b; Zhang et al., 2023a) (Liu et al., 2023b; Shrivastava et al., 2023a; Ding et al., 2023a; Nashid et al., 2024) (Cheng et al., 2024; Tan et al., 2024; Liu et al., 2024b; Deng et al., 2024) (Wu et al., 2024c) - - -  (Pei et al., 2023; Ding et al., 2023b; Zhang et al., 2023a) (Liu et al., 2023b; Ding et al., 2023a; Nashid et al., 2024; Cheng et al., 2024) (Tan et al., 2024; Liu et al., 2024b; Deng et al., 2024; Wu et al., 2024c) -  (Raychev et al., 2016) $\circ$ (Pei et al., 2023; Ding et al., 2023b; Zhang et al., 2023a) (Liu et al., 2023b; Shrivastava et al., 2023a; Ding et al., 2023a; Nashid et al., 2024) (Cheng et al., 2024; Tan et al., 2024; Liu et al., 2024b; Deng et al., 2024) (Wu et al., 2024c) $\circ$ (Pei et al., 2023; Ding et al., 2023b; Zhang et al., 2023a) (Liu et al., 2023b; Shrivastava et al., 2023a; Ding et al., 2023a; Nashid et al., 2024) (Cheng et al., 2024; Tan et al., 2024; Liu et al., 2024b; Deng et al., 2024) (Wu et al., 2024c) $\circ$ (Pei et al., 2023; Ding et al., 2023b; Zhang et al., 2023a) (Liu et al., 2023b; Shrivastava et al., 2023a; Ding et al., 2023a; Nashid et al., 2024) (Cheng et al., 2024; Tan et al., 2024; Liu et al., 2024b; Deng et al., 2024) (Wu et al., 2024c) $\circ$ (Pei et al., 2023; Ding et al., 2023b; Zhang et al., 2023a) (Liu et al., 2023b; Shrivastava et al., 2023a; Ding et al., 2023a; Nashid et al., 2024) (Tan et al., 2024; Liu et al., 2024b; Deng et al., 2024; Wu et al., 2024c) $\circ$ (Pei et al., 2023; Ding et al., 2023b; Zhang et al., 2023a) (Liu et al., 2023b; Shrivastava et al., 2023a; Ding et al., 2023a; Nashid et al., 2024) (Tan et al., 2024; Liu et al., 2024b; Deng et al., 2024; Wu et al., 2024c) - $\circ$ (Wu et al., 2024c)  (Bibaev et al., 2022)
Java  (Liu et al., 2023b; Ding et al., 2023a; Nashid et al., 2024) (Liu et al., 2024b; Deng et al., 2024) - - -  (Liu et al., 2023b; Ding et al., 2023a; Nashid et al., 2024) (Liu et al., 2024b; Deng et al., 2024) - $\circ$ (Liu et al., 2023b; Ding et al., 2023a; Nashid et al., 2024) (Liu et al., 2024b; Deng et al., 2024) $\circ$ (Liu et al., 2023b; Ding et al., 2023a; Nashid et al., 2024) (Liu et al., 2024b; Deng et al., 2024) $\circ$ (Liu et al., 2023b; Ding et al., 2023a; Nashid et al., 2024) (Liu et al., 2024b; Deng et al., 2024) $\circ$ (Liu et al., 2023b; Ding et al., 2023a; Nashid et al., 2024) (Liu et al., 2024b; Deng et al., 2024) $\circ$ (Liu et al., 2023b; Ding et al., 2023a; Nashid et al., 2024) (Liu et al., 2024b; Deng et al., 2024) - - -
Other  (Ding et al., 2023a; Deng et al., 2024; Wu et al., 2024c) (Guan and Treude, 2024) - - -  (Ding et al., 2023a; Deng et al., 2024; Wu et al., 2024c) (Guan and Treude, 2024) - $\circ$ (Ding et al., 2023a; Deng et al., 2024; Wu et al., 2024c) (Guan and Treude, 2024) $\circ$ (Ding et al., 2023a; Deng et al., 2024; Wu et al., 2024c) (Guan and Treude, 2024) $\circ$ (Ding et al., 2023a; Deng et al., 2024; Wu et al., 2024c) (Guan and Treude, 2024) $\circ$ (Ding et al., 2023a; Deng et al., 2024; Wu et al., 2024c) (Guan and Treude, 2024) $\circ$ (Ding et al., 2023a; Deng et al., 2024; Wu et al., 2024c) (Guan and Treude, 2024) - $\circ$ (Wu et al., 2024c; Guan and Treude, 2024) -
@@
Java  (Zhou et al., 2019; Guo et al., 2023a; Wang et al., 2021a) (Wang et al., 2023; Zhou et al., 2024; Bansal et al., 2024) -  (Shahbazi and Fard, 2023) -  (Zhou et al., 2019; Guo et al., 2023a; Wang et al., 2021a) (Wang et al., 2023; Zhou et al., 2024; Bansal et al., 2024) - $\circ$ (Zhou et al., 2019; Guo et al., 2023a; Wang et al., 2021a) (Wang et al., 2023; Zhou et al., 2024; Bansal et al., 2024) $\circ$ (Zhou et al., 2019; Guo et al., 2023a; Wang et al., 2021a) (Wang et al., 2023; Zhou et al., 2024; Bansal et al., 2024) $\circ$ (Zhou et al., 2019; Guo et al., 2023a; Wang et al., 2021a) (Wang et al., 2023; Zhou et al., 2024; Bansal et al., 2024) $\circ$ (Wang et al., 2023; Zhou et al., 2024; Bansal et al., 2024) $\circ$ (Wang et al., 2023; Zhou et al., 2024; Bansal et al., 2024)  (Wang et al., 2021a) - -
@@
Commit Message Generation Python  (Wang et al., 2024f; Eliseeva et al., 2023)  (Wang et al., 2024f; Eliseeva et al., 2023) - - -  (Wang et al., 2024f; Eliseeva et al., 2023) $\circ$ (Wang et al., 2024f; Eliseeva et al., 2023) $\circ$ (Wang et al., 2024f; Eliseeva et al., 2023) $\circ$ (Wang et al., 2024f; Eliseeva et al., 2023) $\circ$ (Wang et al., 2024f; Eliseeva et al., 2023) $\circ$ (Wang et al., 2024f; Eliseeva et al., 2023) - - -
Java  (Xu et al., 2022; Wang et al., 2021b; Vu et al., 2023) (Eliseeva et al., 2023)  (Xu et al., 2022; Wang et al., 2021b) (Vu et al., 2023; Eliseeva et al., 2023) - - -  (Xu et al., 2022; Wang et al., 2021b) (Vu et al., 2023; Eliseeva et al., 2023) $\circ$ (Xu et al., 2022; Wang et al., 2021b; Vu et al., 2023) (Eliseeva et al., 2023) $\circ$ (Xu et al., 2022; Wang et al., 2021b; Vu et al., 2023) (Eliseeva et al., 2023) $\circ$ (Xu et al., 2022; Wang et al., 2021b; Vu et al., 2023) (Eliseeva et al., 2023) $\circ$ (Vu et al., 2023; Eliseeva et al., 2023) $\circ$ (Vu et al., 2023; Eliseeva et al., 2023) - - -
@@
Defect Detection Python  (Wang et al., 2024a) - - -  (Wang et al., 2024a) - $\circ$ (Wang et al., 2024a) $\circ$ (Wang et al., 2024a) $\circ$ (Wang et al., 2024a) $\circ$ (Wang et al., 2024a) $\circ$ (Wang et al., 2024a) - - -
Java  (Li, 2020; Zhang et al., 2023b; Zhou et al., 2012) (Wang et al., 2024a; Xu et al., 2021) - -  (Zhou et al., 2012)  (Li, 2020; Zhang et al., 2023b; Zhou et al., 2012) (Wang et al., 2024a; Xu et al., 2021) - $\circ$ (Li, 2020; Zhang et al., 2023b; Zhou et al., 2012) (Wang et al., 2024a; Xu et al., 2021) $\circ$ (Li, 2020; Zhang et al., 2023b; Zhou et al., 2012) (Wang et al., 2024a; Xu et al., 2021) $\circ$ (Li, 2020; Zhang et al., 2023b; Zhou et al., 2012) (Wang et al., 2024a; Xu et al., 2021) $\circ$ (Li, 2020; Zhang et al., 2023b; Zhou et al., 2012) (Wang et al., 2024a; Xu et al., 2021) $\circ$ (Li, 2020; Zhang et al., 2023b; Zhou et al., 2012) (Wang et al., 2024a; Xu et al., 2021) - - -
Other  (Wang et al., 2024a; Wen et al., 2024b) - - -  (Wang et al., 2024a; Wen et al., 2024b) - $\circ$ (Wang et al., 2024a; Wen et al., 2024b) $\circ$ (Wang et al., 2024a; Wen et al., 2024b) $\circ$ (Wang et al., 2024a; Wen et al., 2024b) $\circ$ (Wang et al., 2024a; Wen et al., 2024b) $\circ$ (Wang et al., 2024a; Wen et al., 2024b) - - -
Program Repair Python  (Chen et al., 2024b; Jimenez et al., 2024)  (Jimenez et al., 2024) -  (Jimenez et al., 2024)  (Chen et al., 2024b; Jimenez et al., 2024)  (Jimenez et al., 2024) $\circ$ (Chen et al., 2024b; Jimenez et al., 2024) $\circ$ (Chen et al., 2024b; Jimenez et al., 2024) $\circ$ (Chen et al., 2024b; Jimenez et al., 2024) $\circ$ (Chen et al., 2024b; Jimenez et al., 2024) $\circ$ (Chen et al., 2024b; Jimenez et al., 2024) - $\circ$ (Jimenez et al., 2024) -
@@
Code Generation (Zan et al., 2022a; Tang et al., 2023; Zhou et al., 2023; Liu et al., 2023c; Nashid et al., 2023; Agashe et al., 2019; Liao et al., 2024; Xie et al., 2024; Li et al., 2024c; Hai et al., 2024; Zheng et al., 2024c; Zhuo et al., 2024) (Zhou et al., 2023; Liu et al., 2023c; Nashid et al., 2023; Lyu et al., 2021) (Zan et al., 2022a; Tang et al., 2023; Zhou et al., 2023; Liu et al., 2023c; Nashid et al., 2023; Agashe et al., 2019; Bi et al., 2024b; Yu et al., 2024; Lyu et al., 2021; Liao et al., 2024; Xie et al., 2024; Li et al., 2024c; Hai et al., 2024; Zheng et al., 2024c; Zhuo et al., 2024)
Code Completion (Agrawal et al., 2024; Pei et al., 2023; Shrivastava et al., 2023b; Zhang et al., 2023a; Liu et al., 2023b; Shrivastava et al., 2023a; Guo et al., 2023b; Cheng et al., 2024; Liu et al., 2024b; Lu et al., 2022) (Agrawal et al., 2024; Shrivastava et al., 2023b; Liu et al., 2023b; Shrivastava et al., 2023a; Guo et al., 2023b; Phan et al., 2024) (Asaduzzaman et al., 2016; Agrawal et al., 2024; Pei et al., 2023; Shrivastava et al., 2023b; Zhang et al., 2023a; Liu et al., 2023b; Shrivastava et al., 2023a; Guo et al., 2023b; Phan et al., 2024; Wu et al., 2024b; Cheng et al., 2024; Wang et al., 2024d; Liu et al., 2024b; Lu et al., 2022)
Code Summarization (Guo et al., 2023a; Haque et al., 2020; Wang et al., 2021a; Bansal et al., 2021, 2023a; Zügner et al., 2021; Lin et al., 2021; Gao et al., 2023; Liu et al., 2021; Guo et al., 2024a; Zhang et al., 2024h; Zhou et al., 2024) (Ding et al., 2023b; Bansal et al., 2021) (Guo et al., 2023a; Ding et al., 2023b; Haque et al., 2020; Bansal et al., 2021, 2023a; Zügner et al., 2021; Lin et al., 2021; Gao et al., 2023; Guo et al., 2024a; Zhang et al., 2024h)
Commit Message Generation (Xu et al., 2022; Wang et al., 2021b; Vu et al., 2023; Wang et al., 2024f, a; Eliseeva et al., 2023) (Vu et al., 2023) (Xu et al., 2022; Nie et al., 2021; Wang et al., 2021b; Vu et al., 2023; Wang et al., 2024f, a)
@@
Defect Detection (Li, 2020; Zhang et al., 2023b; Youm et al., 2015; Zhou et al., 2012; Youm et al., 2017; Xu et al., 2021; Jimenez et al., 2024) (Li, 2020) (Zhang et al., 2023c; Lam et al., 2015; Li, 2020; Zhang et al., 2023b; Youm et al., 2015; Zhou et al., 2012; Zhang et al., 2024a; Youm et al., 2017)
Program Repair (Lutellier et al., 2020; Lin et al., 2022; Nashid et al., 2023; Chen et al., 2021a; Li et al., 2020; Motwani and Brun, 2023) (Lutellier et al., 2020) (Lutellier et al., 2020; Lin et al., 2022; Nashid et al., 2023; Chen et al., 2021a; Kim and Kim, 2019; Li et al., 2020; Jiang et al., 2018; Motwani and Brun, 2023)

### v2

@@ 7.1. Metric
Code Generation Python  (Tang et al., 2023; Agashe et al., 2019; Yu et al., 2024) (Liao et al., 2024; Xie et al., 2024; Li et al., 2024c; Hai et al., 2024) (Zhang et al., 2024a; Zheng et al., 2024b; Zhuo et al., 2024)  (Zheng et al., 2024b)  (Zan et al., 2023, 2022) (Ma et al., 2024b) -  (Tang et al., 2023; Yu et al., 2024; Liao et al., 2024) (Li et al., 2024c; Hai et al., 2024; Zheng et al., 2024b; Zhuo et al., 2024) $\circ$ (Zheng et al., 2024b) $\circ$ (Tang et al., 2023; Agashe et al., 2019; Yu et al., 2024) (Liao et al., 2024; Xie et al., 2024; Li et al., 2024c) (Hai et al., 2024; Zheng et al., 2024b; Zhuo et al., 2024) $\circ$ (Tang et al., 2023; Agashe et al., 2019; Yu et al., 2024) (Liao et al., 2024; Xie et al., 2024; Li et al., 2024c) (Hai et al., 2024; Zheng et al., 2024b; Zhuo et al., 2024) $\circ$ (Tang et al., 2023; Agashe et al., 2019; Yu et al., 2024) (Liao et al., 2024; Xie et al., 2024; Li et al., 2024c) (Hai et al., 2024; Zheng et al., 2024b; Zhuo et al., 2024) $\circ$ (Tang et al., 2023; Yu et al., 2024; Liao et al., 2024) (Li et al., 2024c; Hai et al., 2024; Zheng et al., 2024b; Zhuo et al., 2024) $\circ$ (Tang et al., 2023; Yu et al., 2024; Liao et al., 2024) (Li et al., 2024c; Hai et al., 2024; Zheng et al., 2024b; Zhuo et al., 2024) - $\circ$ (Zan et al., 2023, 2022; Tang et al., 2023) (Yu et al., 2024; Xie et al., 2024; Li et al., 2024c; Hai et al., 2024) (Zhang et al., 2024a; Zheng et al., 2024b; Zhuo et al., 2024) -
@@
Code Completion Python  (Pei et al., 2023; Ding et al., 2023b; Zhang et al., 2023a) (Liu et al., 2023c; Shrivastava et al., 2023a; Ding et al., 2023a; Nashid et al., 2024) (Cheng et al., 2024; Tan et al., 2024; Liu et al., 2024b; Deng et al., 2024) (Wu et al., 2024b) - - -  (Pei et al., 2023; Ding et al., 2023b; Zhang et al., 2023a) (Liu et al., 2023c; Ding et al., 2023a; Nashid et al., 2024; Cheng et al., 2024) (Tan et al., 2024; Liu et al., 2024b; Deng et al., 2024; Wu et al., 2024b) -  (Raychev et al., 2016) $\circ$ (Pei et al., 2023; Ding et al., 2023b; Zhang et al., 2023a) (Liu et al., 2023c; Shrivastava et al., 2023a; Ding et al., 2023a; Nashid et al., 2024) (Cheng et al., 2024; Tan et al., 2024; Liu et al., 2024b; Deng et al., 2024) (Wu et al., 2024b) $\circ$ (Pei et al., 2023; Ding et al., 2023b; Zhang et al., 2023a) (Liu et al., 2023c; Shrivastava et al., 2023a; Ding et al., 2023a; Nashid et al., 2024) (Cheng et al., 2024; Tan et al., 2024; Liu et al., 2024b; Deng et al., 2024) (Wu et al., 2024b) $\circ$ (Pei et al., 2023; Ding et al., 2023b; Zhang et al., 2023a) (Liu et al., 2023c; Shrivastava et al., 2023a; Ding et al., 2023a; Nashid et al., 2024) (Cheng et al., 2024; Tan et al., 2024; Liu et al., 2024b; Deng et al., 2024) (Wu et al., 2024b) $\circ$ (Pei et al., 2023; Ding et al., 2023b; Zhang et al., 2023a) (Liu et al., 2023c; Shrivastava et al., 2023a; Ding et al., 2023a; Nashid et al., 2024) (Tan et al., 2024; Liu et al., 2024b; Deng et al., 2024; Wu et al., 2024b) $\circ$ (Pei et al., 2023; Ding et al., 2023b; Zhang et al., 2023a) (Liu et al., 2023c; Shrivastava et al., 2023a; Ding et al., 2023a; Nashid et al., 2024) (Tan et al., 2024; Liu et al., 2024b; Deng et al., 2024; Wu et al., 2024b) - $\circ$ (Wu et al., 2024b)  (Bibaev et al., 2022)
Java  (Liu et al., 2023c; Ding et al., 2023a; Nashid et al., 2024) (Liu et al., 2024b; Deng et al., 2024) - - -  (Liu et al., 2023c; Ding et al., 2023a; Nashid et al., 2024) (Liu et al., 2024b; Deng et al., 2024) - $\circ$ (Liu et al., 2023c; Ding et al., 2023a; Nashid et al., 2024) (Liu et al., 2024b; Deng et al., 2024) $\circ$ (Liu et al., 2023c; Ding et al., 2023a; Nashid et al., 2024) (Liu et al., 2024b; Deng et al., 2024) $\circ$ (Liu et al., 2023c; Ding et al., 2023a; Nashid et al., 2024) (Liu et al., 2024b; Deng et al., 2024) $\circ$ (Liu et al., 2023c; Ding et al., 2023a; Nashid et al., 2024) (Liu et al., 2024b; Deng et al., 2024) $\circ$ (Liu et al., 2023c; Ding et al., 2023a; Nashid et al., 2024) (Liu et al., 2024b; Deng et al., 2024) - - -
Other  (Ding et al., 2023a; Deng et al., 2024; Wu et al., 2024b) (Guan and Treude, 2024) - - -  (Ding et al., 2023a; Deng et al., 2024; Wu et al., 2024b) (Guan and Treude, 2024) - $\circ$ (Ding et al., 2023a; Deng et al., 2024; Wu et al., 2024b) (Guan and Treude, 2024) $\circ$ (Ding et al., 2023a; Deng et al., 2024; Wu et al., 2024b) (Guan and Treude, 2024) $\circ$ (Ding et al., 2023a; Deng et al., 2024; Wu et al., 2024b) (Guan and Treude, 2024) $\circ$ (Ding et al., 2023a; Deng et al., 2024; Wu et al., 2024b) (Guan and Treude, 2024) $\circ$ (Ding et al., 2023a; Deng et al., 2024; Wu et al., 2024b) (Guan and Treude, 2024) - $\circ$ (Wu et al., 2024b; Guan and Treude, 2024) -
@@
Java  (Zhou et al., 2019; Guo et al., 2023b; Wang et al., 2021b) (Wang et al., 2023; Zhou et al., 2024; Bansal et al., 2024) -  (Shahbazi and Fard, 2023) -  (Zhou et al., 2019; Guo et al., 2023b; Wang et al., 2021b) (Wang et al., 2023; Zhou et al., 2024; Bansal et al., 2024) - $\circ$ (Zhou et al., 2019; Guo et al., 2023b; Wang et al., 2021b) (Wang et al., 2023; Zhou et al., 2024; Bansal et al., 2024) $\circ$ (Zhou et al., 2019; Guo et al., 2023b; Wang et al., 2021b) (Wang et al., 2023; Zhou et al., 2024; Bansal et al., 2024) $\circ$ (Zhou et al., 2019; Guo et al., 2023b; Wang et al., 2021b) (Wang et al., 2023; Zhou et al., 2024; Bansal et al., 2024) $\circ$ (Wang et al., 2023; Zhou et al., 2024; Bansal et al., 2024) $\circ$ (Wang et al., 2023; Zhou et al., 2024; Bansal et al., 2024)  (Wang et al., 2021b) - -
@@
Commit Message Generation Python  (Wang et al., 2024b; Eliseeva et al., 2023)  (Wang et al., 2024b; Eliseeva et al., 2023) - - -  (Wang et al., 2024b; Eliseeva et al., 2023) $\circ$ (Wang et al., 2024b; Eliseeva et al., 2023) $\circ$ (Wang et al., 2024b; Eliseeva et al., 2023) $\circ$ (Wang et al., 2024b; Eliseeva et al., 2023) $\circ$ (Wang et al., 2024b; Eliseeva et al., 2023) $\circ$ (Wang et al., 2024b; Eliseeva et al., 2023) - - -
Java  (Xu et al., 2022; Wang et al., 2021a; Vu et al., 2023) (Eliseeva et al., 2023)  (Xu et al., 2022; Wang et al., 2021a) (Vu et al., 2023; Eliseeva et al., 2023) - - -  (Xu et al., 2022; Wang et al., 2021a) (Vu et al., 2023; Eliseeva et al., 2023) $\circ$ (Xu et al., 2022; Wang et al., 2021a; Vu et al., 2023) (Eliseeva et al., 2023) $\circ$ (Xu et al., 2022; Wang et al., 2021a; Vu et al., 2023) (Eliseeva et al., 2023) $\circ$ (Xu et al., 2022; Wang et al., 2021a; Vu et al., 2023) (Eliseeva et al., 2023) $\circ$ (Vu et al., 2023; Eliseeva et al., 2023) $\circ$ (Vu et al., 2023; Eliseeva et al., 2023) - - -
@@
Defect Detection Python  (Wang et al., 2024c) - - -  (Wang et al., 2024c) - $\circ$ (Wang et al., 2024c) $\circ$ (Wang et al., 2024c) $\circ$ (Wang et al., 2024c) $\circ$ (Wang et al., 2024c) $\circ$ (Wang et al., 2024c) - - -
Java  (Li, 2020; Zhang et al., 2023b; Zhou et al., 2012) (Wang et al., 2024c; Xu et al., 2021) - -  (Zhou et al., 2012)  (Li, 2020; Zhang et al., 2023b; Zhou et al., 2012) (Wang et al., 2024c; Xu et al., 2021) - $\circ$ (Li, 2020; Zhang et al., 2023b; Zhou et al., 2012) (Wang et al., 2024c; Xu et al., 2021) $\circ$ (Li, 2020; Zhang et al., 2023b; Zhou et al., 2012) (Wang et al., 2024c; Xu et al., 2021) $\circ$ (Li, 2020; Zhang et al., 2023b; Zhou et al., 2012) (Wang et al., 2024c; Xu et al., 2021) $\circ$ (Li, 2020; Zhang et al., 2023b; Zhou et al., 2012) (Wang et al., 2024c; Xu et al., 2021) $\circ$ (Li, 2020; Zhang et al., 2023b; Zhou et al., 2012) (Wang et al., 2024c; Xu et al., 2021) - - -
Other  (Wang et al., 2024c; Wen et al., 2024) - - -  (Wang et al., 2024c; Wen et al., 2024) - $\circ$ (Wang et al., 2024c; Wen et al., 2024) $\circ$ (Wang et al., 2024c; Wen et al., 2024) $\circ$ (Wang et al., 2024c; Wen et al., 2024) $\circ$ (Wang et al., 2024c; Wen et al., 2024) $\circ$ (Wang et al., 2024c; Wen et al., 2024) - - -
Program Repair Python  (Chen et al., 2024; Jimenez et al., 2024)  (Jimenez et al., 2024) -  (Jimenez et al., 2024)  (Chen et al., 2024; Jimenez et al., 2024)  (Jimenez et al., 2024) $\circ$ (Chen et al., 2024; Jimenez et al., 2024) $\circ$ (Chen et al., 2024; Jimenez et al., 2024) $\circ$ (Chen et al., 2024; Jimenez et al., 2024) $\circ$ (Chen et al., 2024; Jimenez et al., 2024) $\circ$ (Chen et al., 2024; Jimenez et al., 2024) - $\circ$ (Jimenez et al., 2024) -
@@
Code Generation (Zan et al., 2022; Tang et al., 2023; Zhou et al., 2023; Liu et al., 2023b; Nashid et al., 2023; Agashe et al., 2019; Liao et al., 2024; Xie et al., 2024; Li et al., 2024c; Hai et al., 2024; Zheng et al., 2024b; Zhuo et al., 2024) (Zhou et al., 2023; Liu et al., 2023b; Nashid et al., 2023; Lyu et al., 2021) (Zan et al., 2022; Tang et al., 2023; Zhou et al., 2023; Liu et al., 2023b; Nashid et al., 2023; Agashe et al., 2019; Bi et al., 2024; Yu et al., 2024; Lyu et al., 2021; Liao et al., 2024; Xie et al., 2024; Li et al., 2024c; Hai et al., 2024; Zheng et al., 2024b; Zhuo et al., 2024)
Code Completion (Agrawal et al., 2024; Pei et al., 2023; Shrivastava et al., 2023b; Zhang et al., 2023a; Liu et al., 2023c; Shrivastava et al., 2023a; Guo et al., 2023a; Cheng et al., 2024; Liu et al., 2024b; Lu et al., 2022) (Agrawal et al., 2024; Shrivastava et al., 2023b; Liu et al., 2023c; Shrivastava et al., 2023a; Guo et al., 2023a; Phan et al., 2024) (Asaduzzaman et al., 2016; Agrawal et al., 2024; Pei et al., 2023; Shrivastava et al., 2023b; Zhang et al., 2023a; Liu et al., 2023c; Shrivastava et al., 2023a; Guo et al., 2023a; Phan et al., 2024; Wu et al., 2024a; Cheng et al., 2024; Wang et al., 2024g; Liu et al., 2024b; Lu et al., 2022)
Code Summarization (Guo et al., 2023b; Haque et al., 2020; Wang et al., 2021b; Bansal et al., 2021, 2023a; Zügner et al., 2021; Lin et al., 2021; Gao et al., 2023; Liu et al., 2021; Guo et al., 2024b; Zhang et al., 2024b; Zhou et al., 2024) (Ding et al., 2023b; Bansal et al., 2021) (Guo et al., 2023b; Ding et al., 2023b; Haque et al., 2020; Bansal et al., 2021, 2023a; Zügner et al., 2021; Lin et al., 2021; Gao et al., 2023; Guo et al., 2024b; Zhang et al., 2024b)
Commit Message Generation (Xu et al., 2022; Wang et al., 2021a; Vu et al., 2023; Wang et al., 2024b, c; Eliseeva et al., 2023) (Vu et al., 2023) (Xu et al., 2022; Nie et al., 2021; Wang et al., 2021a; Vu et al., 2023; Wang et al., 2024b, c)
@@
Defect Detection (Li, 2020; Zhang et al., 2023b; Youm et al., 2015; Zhou et al., 2012; Youm et al., 2017; Xu et al., 2021; Jimenez et al., 2024) (Li, 2020) (Zhang et al., 2023d; Lam et al., 2015; Li, 2020; Zhang et al., 2023b; Youm et al., 2015; Zhou et al., 2012; Zhang et al., 2024d; Youm et al., 2017)
Program Repair (Lutellier et al., 2020; Lin et al., 2022; Nashid et al., 2023; Chen et al., 2021b; Li et al., 2020; Motwani and Brun, 2023) (Lutellier et al., 2020) (Lutellier et al., 2020; Lin et al., 2022; Nashid et al., 2023; Chen et al., 2021b; Kim and Kim, 2019; Li et al., 2020; Jiang et al., 2018; Motwani and Brun, 2023)